OpenStudy (anonymous):

The average of four numbers is 27. If the first number is increased by k, the second by (k + 7), the third by (2k + 3) and the fourth is decreased by 2, the new average of the four numbers becomes 32. Find the value of k.

OpenStudy (anonymous):

27*4=108,k+k+7+2k+3+k-2=32,k=4.8

OpenStudy (anonymous):

(x + y + a + b)/4 = 27 (x+k+y+k+7+a+2k+3+b-2)/4 = 32 By solving k=3